one mole of gas allowed to expand reversibly i.e. against confining pressure that is at all times infinitesimally less than the gas pressure, from an initial pressure of 10atm to a final pressure of 0.4atm, the temperature being kept constant at 0oC.

a. how much work is done by the gas?

b. what is the change in initial energy and enthalpy?

c. How much heat is absorbed?


1
Expert's answer
2021-06-25T05:41:37-0400

a) work done= P1/T1-P2/T2

= 10/273 -0.4/273

= 0.0366 -0.0015

= 0.0351J

b) ∆H = m x s x ∆T 

= 4.021 × 273

= 1097.7J

c) Heat absorbed = 273 × 4.021 × -9.6

= 10528.24J